I need to know the formula for this. Thanks!
Ann drove 60 kph and then 80 kph. The total distance is 540 km and took 8 hours, how far did she drive during 60 kph?

Answer by mananth(16946)

First Part x km
Second Part -x 540 km
Total time 8 hours
First Part time 60 / x
Second Part time 80 / (- 540 - x)

Time first part + time second part = 8 hours

x/ 60 +( 540 -x)/ 80 = 8

multiply the equation by the LCD
we get

80 x+ 60 ( 540 -x)= 8 * 60 * 80

80 x+ 32400 - 60 x = 38400
20 x+ 32400 = 38400
/ 20
4 x 1620 x+ = 1920
4 x = 1920 - 1620
4 x = 300
x= 75 km First part